OpenStudy (gabylovesyou):

Using the pythagorean identity, prove (sin 0) / sqrt(1-sin^2 0) = tan 0 In which quadrants is this equation true ?

OpenStudy (gabylovesyou):

@phi

OpenStudy (gabylovesyou):

@jigglypuff314

OpenStudy (phi):

I assume you mean \( \theta\) and not 0, right ? just use x for the angle.

OpenStudy (gabylovesyou):

yes.

OpenStudy (phi):

\[ \frac{\sin x }{\sqrt{1- \cos^2 x} }= \tan x \] ?

OpenStudy (gabylovesyou):

sin*

OpenStudy (phi):

\[ \frac{\sin x }{\sqrt{1- \sin^2 x} }= \tan x \] by pythagorean identify, I assume they mean use \[ \sin^2 x + \cos^2 x= 1 \] to simplify the bottom square root stuff.

OpenStudy (gabylovesyou):

yes.

OpenStudy (phi):

can you simplify the bottom part ?

OpenStudy (gabylovesyou):

idk how to

OpenStudy (phi):

in sin^2 x + cos^2 x = 1 add -sin^2 x to both sides and simplify \[ \sin^2 x - \sin^2 x+ \cos^2 x = 1-\sin^2 x \\\cos^2 x = 1-\sin^2 x\]

OpenStudy (gabylovesyou):

ok and then

OpenStudy (phi):

do you see that you can replace 1- sin^2 x with cos^2 x ?

OpenStudy (gabylovesyou):

ohh yes

OpenStudy (phi):

what do you get if you "swap out" 1-sin^2 x with cos^2 x ?

OpenStudy (gabylovesyou):

(sin theta) / sqrt (cos^2 theta) = tan theta

OpenStudy (phi):

ok, and what is the square root of a square ?

OpenStudy (gabylovesyou):

im so lost. what do you mean ?

OpenStudy (phi):

if you had \[ \sqrt{x^2} = x\] we use that same idea for sqr(cos^2 x)

OpenStudy (gabylovesyou):

oh ok.

OpenStudy (phi):

in other words \[ \sqrt{\cos^2 x} = \cos x \]

OpenStudy (gabylovesyou):

ok

OpenStudy (phi):

what do you have now ?

OpenStudy (gabylovesyou):

sin theta / cos theta = tan theta

OpenStudy (phi):

yes. but (unfortunately) there is a small detail. we assume we take the positive value of the square root. we should write the bottom as | cos x | (absolute value ) to show this on the right side, tan x = sin x / cos x (by definition) so you have \[ \frac{\sin x}{ | \cos x | } = \frac{\sin x }{\cos x} \] those will be identical as long as cos x is positive (when cos x is negative, the left side with the absolute values will make it positive, and different from the right side)

OpenStudy (phi):

so that identity is true when cos is positive. any idea what quadrants where cos is positive ?

OpenStudy (gabylovesyou):

ohhh

OpenStudy (phi):

I remember that "cos" goes with "x"